### v2

# Reconstructing Sepsis Trajectories from Clinical Case Reports using LLMs: the Textual Time Series Corpus for Sepsis

Abstract
Clinical case reports and discharge summaries may be the most complete and accurate summarization of patient encounters, yet they are finalized, i.e., timestamped after the encounter. Complementary data structured streams become available sooner but suffer from incompleteness. To train models and algorithms on more complete and temporally fine-grained data, we construct a pipeline to phenotype, extract, and annotate time-localized findings within case reports using large language models. We apply our pipeline to generate an open-access textual time series corpus for Sepsis-3 comprising 2,139 case reports from the Pubmed-Open Access (PMOA) Subset. To validate our system, we apply it on PMOA and timeline annotations from I2B2/MIMIC-IV and compare the results to physician-expert annotations. We show high recovery rates of clinical findings (event match rates: O1-preview–0.755, Llama 3.3 70B Instruct–0.753) and strong temporal ordering (concordance: O1-preview–0.932, Llama 3.3 70B Instruct–0.932). Our work characterizes the ability of LLMs to time-localize clinical findings in text, illustrating the limitations of LLM use for temporal reconstruction and providing several potential avenues of improvement via multimodal integration.
keywords:
Clinical Time Series, Sepsis
1 Introduction
The Third International Consensus Definitions for Sepsis (Sepsis-3) are the benchmark definitions for sepsis, the dysfunctional immune response to an infection, which are used in trial eligibility criteria [1] and phenotyping studies [2] . In the critical care literature, sepsis is seen as a heterogeneous disease, a pathway arrived upon from a multitude of infectious origins, anatomical sites, and comorbid profiles. There is increasing recognition of the importance of time in understanding sepsis and sepsis progression, as seen in early-warning systems [3] , variation in prediction utility [4] , and endotype characterization [5] .
Meanwhile, structured data streams, which are the data resources these tools use, are often incomplete records with insufficient information to render or confirm a diagnosis of sepsis [6]. A more complete alternative is the discharge summary, or more generally, the case report, in which a clinician documents the clinical findings pertinent to the case and care management. The impracticality of using the case report is that it is only written in full after the encounter is over, i.e., with a less granular timestamp than the events contained within. In principle, having a more complete, fine-grained temporal record could enable better temporal analyses of patients with sepsis, leading to better understanding and treatment improvements.
At present, the (semi-)public critical care repositories at the research frontier contain high resolution structured signals and low resolution textual signals. While several concerted efforts have focused on extracting temporal information, they have largely focused on temporal relations between clinical concepts [7] rather than on event timing. Those that have focused on timing [8, 9] used excerpted summaries and relatively small sample sizes.
In order to train sepsis models that can consider sources of information references in free-text and that have fine time-granularity, we seek to generate larger samples of textual time series. To do so, we introduce a novel Sepsis-3 textual time series corpus from the Pubmed Open Access (PMOA) Subset. We adopt the LLM-as-annotator approach, evaluating its ability to extract clinical findings and its ability to accurately associate them with timestamps. We augment existing clinical annotations with our own on PMOA, and contrast LLM annotations with our own and that of Frattallone et al. (2024) [9] . To provide checks on their performance at scale, we consider a larger subsample for inter-LLM annotator agreement, and show the performance levels across several model sizes, which provides valuable context for cost-based analyses. Finally, these assessments provide evidence of the LLM annotations quality and our pipeline, which are then applied to generate the textual time series corpus for sepsis (T2S2).
1.1 Related work
Numerous critical care and sepsis datasets are available (semi-)publicly [10, 11] , including several with de-identified clinical discharge summaries: MIMIC-III [12] and MIMIC-IV [13] . Sepsis phenotyping (including Sepsis-3) has been conducted using a variety of methods on structured data [14] , although structured data is often highly missing, leading to suboptimal phenotypes [15]. Additionally, the timing of sepsis phenotyping is crucial, in that delays in sepsis prediction can nullify predictiveness and lead to burdensome and unhelpful analytic models [4] .
Several works have used the i2b2 (a competition subset of MIMIC-II/III) to construct timelines of clinical concepts [8, 9] . In Leeuwenberg et al. [8] , the event spans are already given, and they focus on an excerpt of the discharge summary. Hence, we focus our direct comparison on that of Frattallone et al [9] . In contrast to these works, we adopt an annotation process focused on temporal assignment to clinical finding rather than i2b2 clinical concept [16] , which allows for greater specificity of the event in the absence of additional context. Additionally, as compared to Frattallone et al. [9] , we focus on using text alone for clinical time series construction, which is necessary for PMOA since no complementary data source is available.
Broadly in the medical domain, LLMs are showing promise in medical tasks such as summarization [17] . In fact, some findings suggest that medical fine-tuning rarely leads to improved model over their corresponding foundation models [18] . In our analysis, we primarily use foundation LLMs and show that their scale and instruction-tuning are central to their performance.
1.2 Contributions
Our work highlights the following contributions: (1) we present the first textual time series corpus for sepsis, comprising 2,139 open-access case reports identified as having sepsis; (2) we develop an analytic pipeline to identify Sepsis-3 from clinical text and to extract time-associated clinical findings; (3) we annotate time-associated clinical findings on a random subset of the case reports; (4) we develop and adapt evaluation criteria for textual time series, where matched clinical findings are not expected to be identical text strings; and, (5) in multiple datasets, we validate the performance of our pipeline, showing strong annotation performance compared to manual annotation and improved temporal results compared to existing methods.
2 Methods
We define the task specification, relevant terms, the extraction and annotation procedure, and the assessments.
i2m4
[…]
Yes 2 0
No 0 18
L3.1 8B Yes No
Yes 2 7
No 0 11
sepsis-100
[…]
Yes 78 7
No 8 7
L3.1 8B Yes No
Yes 81 12
No 5 2
Figure 1: PMOA T2S2 pipeline (left) and Sepsis-3 confusion matrices (right)
Our task is the following: given a biomedical free-text corpus, construct a textual time series case report corpus of patients with Sepsis-3. We define a textual time series as a list of clinical findings, each with a timestamp (absolute or relative to case presentation) for an individual. A clinical finding is a free-text specification of an entity pertaining to or with the potential to affect the person’s health.
Compared to clinical concept annotations [16, 7] , we choose to expand the text span determination to increase the specificity of the meaning attached to the finding. We note the following distinctions of our notion of clinical finding with that of i2b2 clinical concept, as per the i2b2 guideline specifications [16] :
  1. 1.
    Clinical findings are not limited to one prepositional phrase following a markable concept (e.g., “pain in chest that radiates substernally”). Instead of splitting into concepts “pain in chest” and possibly “substernal radiation”, where the former lacks specificity and the latter lacks origin, a single finding is preferred.
  2. 2.
    Related conjunctions should be split into component findings. Instead of “metastases in the liver and pancreas”, we prefer two more specific clinical findings “metastasis in the liver” and “metastasis in the pancreas”.
The other design choice we make is to alter the time expression to disambiguate timing by capturing the start time of the perception of the clinical event. For example, if a patient has a “three-day history of fever”, the clinical event “fever” is assigned to -72 hours, since it is first perceived then. If the patient has a “history of smoking”, the start time is fully ambiguous, so the time of perception is the time that smoking is assessed, e.g., at the start of the encounter, and the event is “history of smoking”, rather than “smoking”. Our clinical finding is defined as a contiguous text span representing a temporally localizable patient event, adjusted for conjunction splitting and temporal perception.
Given this specification, our subtasks were: case report detection, Sepsis-3 detection, clinical finding extraction, and timestamp determination. We constructed a pipeline for these steps and their assessment (Figure 1 (left), showing the process for the PMOA dataset, with an analogous process used for i2m4).
2.1 Data
We used two data sources: (1) the absolute timeline annotations [9] , and (2) the Pubmed Open Access subset [19] (see Table 2.1.1).
2.1.1 i2m4 Dataset
The absolute timeline annotations [9] came from physician annotations of 20 discharge summaries following the i2b2 concept and annotation guidelines (note, the original competition annotation focused on discharge summary excerpts rather the entire note), with the addition of interval timestamping of the identified concepts. For our analysis, we used the lower bound of the interval (specified as a mean time) as it denotes the time when the finding may first be observed. We used only the probabilistic annotations, since the bounded annotation lower bound could precede the actual occurrence of the finding. Fifteen of the summaries come from i2b2 and five are from MIMIC IV (annotated similarly)—hence we call this dataset “i2m4”.
[…]
Dataset summary. Bold cells indicate contributed labeling Corpus i2m4 Sepsis-10   Sepsis-100 N (unlabeled) 310; 331,794 2,139 2,139 N (labeled) 20 10 100 Source i2b2; MIMIC IV PMOA PMOA Note type discharge summary case report case report Labeler clinicians clinicians O1-preview Phenotyper clinicians clinicians QwQ-32B-Preview Methods O1-preview O1-preview DeepSeek R1 IQ1 DeepSeek R1 IQ1 DeepSeek R1 IQ1 L3.3 70B Instruct L3.3 70B Instruct L3.3 70B Instruct L3.1 70B Instruct L3.1 70B Instruct L3.1 70B Instruct Mixtral 8x7B Mixtral 8x7B Mixtral 8x7B  Instruct v0.1  Instruct v0.1  Instruct v0.1 L3.1 8B Instruct L3.1 8B Instruct L3.1 8B Instruct Ablations No role No role No role Base: No conjunction No conjunction No conjunction L3.3 70B Instruct  instruction  instruction  instruction No few-shot prompting No few-shot prompting No few-shot prompting Sensitivity analyses With upper bound With upper bound With upper bound With upper bound + With upper bound + With upper bound +  i2b2 event typing  i2b2 event typing  i2b2 event typing
For Sepsis-3 determination, we had a clinician review the discharge summaries and assign each case as meeting the Sepsis-3 definition. To develop a computational approach, we constructed a prompt specific to Sepsis-3, which describes the qSOFA procedure for determining a Sepsis-3 diagnosis, and used it as query to Llama-3.1-8B-Instruct (L3.1 8B) and QwQ-32B-Preview (Qwen). We used the phenotyping prompt:
[…]
We constructed confusion matrices to contrast the Sepsis-3 phenotyping abilities of the LLM-query against clinician determinations (Figure 1, top right).
2.1.2 PMOA (sepsis-10, sepsis-100) Dataset
The second dataset comes from the Pubmed Open Access Subset, an open-access, free-text corpus of roughly 1.5 million biomedical publications [19] . Because the majority of these publications are not case reports, we built an extraction pipeline to recover critical care case reports, i.e., a clinical research informatics critical care cohort.
We used the PMOA Subset (version as of April 25, 2024). We removed the header and footer of the free-text which contained the Title, Abstract, and References sections, and retained the ==== Body section. To identify potential case reports, we required that the free-text match the case-insensitive regular expressions case (report|presenta) and year-? ?old. We compared the query output to PubMed case report metadata and found our regex approach had much higher specificity in a sample of 20 random reports.
To identify potential sepsis case reports, we filtered the potential case reports using the case-insensitive regular expressions (sepsi||septic) and (critical||intensive) care. Then, we used the an LLM query to extract sepsis cases, as well as the number of cases present, and the age and gender. We used GPT-4o and Llama-3.1-8B-Instruct for the phenotyping query screen. Reports were omitted if the number of cases identified was not equal to 1. As a screen, we chose to include the case report if either GPT-4o or Llama-3.1-8B-Instruct identified the report as having Sepsis-3.
This process resulted in 2,139 identified Sepsis-3 case reports, respectively, among which 10 and 100 were sampled for further review. The sepsis-10 sample was annotated by a clinical expert to provide a “ground truth” labeling, and the sepsis-100 sample was used as a larger sample for testing. In sepsis-100, to better assess Sepsis-3 presence, we used the QwQ-32B-Preview model and Sepsis-3 phenotyping prompt above (since it had strong performance characteristics in the i2m4 dataset) to validate the GPT-4o and Llama-3.1-8B-Instruct responses (Figure 1, bottom right).
2.2 Textual Time Series Annotations
For our three datasets—i2m4, sepsis-10, and sepsis-100—we used the following LLMs: O1-preview (o1-preview-2024-09-12), DeepSeek-R1-UD-IQ1_S, Llama-3.3-70B-Instruct, Llama-3.1-70B-Instruct, Mixtral-8x7B-Instruct v0.1, and Llama-3.1-8B-Instruct. For data privacy reasons, we did not submit any i2m4 data queries to O1-preview and ran the other models locally. For the sepsis-100 dataset, we did not have access to clinician annotations, so our inter-LLM annotator assessments used the O1-preview annotations as reference. We attempted to use QwQ-32B-Preview and Qwen-2.5-72B-Instruct for clinical finding annotations but were unable to reliably extract query results. We adapted the Llama query using the system/user/assistant query template. The main annotation prompt used is the following:
You are a physician. Extract the clinical events and the related time stamp from the case report. The admission event has timestamp 0. If the event is not available, we treat the event, e.g. current main clinical diagnosis or treatment with timestamp 0. The events happened before event with 0 timestamp have negative time, the ones after the event with 0 timestamp have positive time. The timestamp are in hours. The unit will be omitted when output the result. If there is no temporal information of the event, please use your knowledge and events with temporal expression before and after the events to provide an approximation. We want to predict the future events given the events happened in history. For example, here is the case report.
An 18-year-old male was admitted to the hospital with a 3-day history of fever and rash. Four weeks ago, he was diagnosed with acne and received the treatment with minocycline, 100 mg daily, for 3 weeks. With increased WBC count, eosinophilia, and systemic involvement, this patient was diagnosed with DRESS syndrome. The fever and rash persisted through admission, and diffuse erythematous or maculopapular eruption with pruritus was present. One day later the patient was discharged.
Let’s find the locations of event in the case report, it shows that four weeks ago of fever and rash, four weeks ago, he was diagnosed with acne and receive treatment. So the event of fever and rash happen four weeks ago, 672 hours, it is before admitted to the hospital, so the time stamp is -672. diffuse erythematous or maculopapular eruption with pruritus was documented on the admission exam, so the time stamp is 0 hours, since it happens right at admission. DRESS syndrome has no specific time, but it should happen soon after admission to the hospital, so we use our clinical judgment to give the diagnosis of DRESS syndrome the timestamp 0. Then the output should look like
18 years old $|$ 0
male $|$ 0
admitted to the hospital $|$ 0
fever $|$ -72
rash $|$ -72
acne $|$ -672
minocycline $|$ -672
increased WBC count $|$ 0
eosinophilia|| 0
systemic involvement|| 0
diffuse erythematous or maculopapular eruption|| 0
pruritis $|$ 0
DRESS syndrome $|$ 0
fever persisted $|$ 0
rash persisted $|$ 0
discharged $|$ 24
[…]
For sensitivity analysis to the prompt used, we omitted the color-coded parts in turn to give us the following ablations: no-role playing (no-role), zero-shot prompting (zero-shot), and no conjunction expansion (no expansion). We also tested with augmentations, where we add to the query a request for the time interval (Interval) or time interval and i2b2 event type (Interval+Type), where event type is one of: Factual, Possible, Hypothetical, Conditional, Negated, Historical, Uncertain. For the augmentations, both the original prompt’s instructions and few-shot examples were modified accordingly.
For the i2m4 dataset, we also compared against a BERT-based baseline (BERT 5-CV), our adaption of the Frattallone et al. model (2024) [9] . In their work, they transformed the temporal prediction task into multiclass classification and used a multilayer perceptron (MLP) head with a softmax and cross-entropy loss for training. To predict relative time (hours from case presentation), we made the following adaptations. Similarly to the classifier module, we used a 3-layer MLP regressor head from the output of the BERT event span features and output the prediction means and standard deviations. The regression task (predicting the lower bound) was optimized using L1 loss. We added the BERT features SUBJ_START and SUBJ_END tokens surrounding the event span [20] and removed the dropout layers inside the MLP heads, leaving only one at the final layer. Regarding the timestamps used for prediction, while many events occurred near the admission time, a subset extended months or years away, complicating the design of a regressor capable of both spanning wide magnitudes and maintaining precision for the majority of the means close to the admission time. To address this, we filtered the findings by their magnitudes, retaining them if the means were within two weeks of the admission time and the standard deviations were under two weeks. Filtering reduced the dataset size (in number of findings) from 3,472 to 3,350. Training was carried out for 50 epochs with batch size 32 and learning rate 3e-5. The model training used 5-fold cross-validation.
2.3 Assessments
We assessed the textual time series with the following methods: event match rate, concordance (c-index), median absolute time discrepancy, and log-time discrepancy. To determine event matches between two lists of strings, one for the label clinical findings and one for the predicted clinical findings, we used a recursive best match procedure, where a match is identified by the distance between two strings (see Appendix 2 for pseudocode). Specifically, the recursive best match was used to reduce erroneous surjective matches, i.e., multiple events from the predicted event list mapping to the same event in the reference list. For each reference event, if multiple predicted events matched, the lowest distance match was selected (with event ordering used to break distance ties), and both the predicted and referent events were removed from the lists for the recursive call to match events. Then all matched events from the recursive calls were returned. To compute distance, we tried (1) Levenshtein distance, (2) cosine distance between mean sentence embeddings of BERT (bert-base-uncased) and (3) cosine distance between mean sentence embeddings of PubMedBert (S-PubMedBert-MS-MARCO) [21]. Upon manual review of event matches, we selected option 3 with a cosine similarity cutoff of 0.1.
Among event matches, we then considered the time-ordering and time-discrepancy between label timestamps and predicted timestamps. We computed the concordance, or c-index, which corresponds to the probability of correctly time-ordering a random pair of predicted findings. The median of absolute error of time discrepancies was used to avoid the outsize impact of large discrepancies that would be captured in the mean absolute error. For computing log-time discrepancies, we tranformed the non-negative discrepancies with the function: $\log(1+\Delta t)$, where $\Delta t$ is the time discrepancy (absolute) per clinical finding in hours. Finding-level cumulative time discrepancy plots order the discrepancies from small to large to show the fraction of discrepancies being smaller than some discrepancy threshold. Then, the area under the log-time curve (AULTC) is the area under the log-time curve of the cumulative time discrepancy plot, which we describe next.
2.3.1 Log-Time Cumulative Distribution Function
Given we have $k$ event matches combined across case reports, we want to compute a measure of the discrepancy between the predicted times $t^{p}_{i}$ and the manual reference times $t^{r}_{i}$ for $i={1,\dots,k}$ that captures the quality of time recovery across time scales.
We define the log-time cumulative distribution function as follows. For each matched event $i$, compute the discrepancy $s_{i}=|t^{p}_{i}-t^{r}_{i}|$. then truncate at the maximum cutoff $S_{\text{max}}$ (e.g., 1 year = 8760 hours): $s^{\prime}_{i}=\min(s_{i},S_{\text{max}})$.
Apply the log transformation: $x_{i}=\log(1+s^{\prime}_{i})$ Sort $x_{i}$ in ascending order: $x_{(1)}\leq x_{(2)}\leq\cdots\leq x_{(k)}$. The CDF $F(x)$ is defined as the fraction of log-transformed discrepancies less than or equal to $x$: $F(x)=\frac{1}{k}\sum_{i=1}^{k}\mathbf{1}_{\{x_{(i)}\leq x\}}$ where $\mathbf{1}$ is the indicator function. Then, we get the following boundary values: $F(0)\geq 0$ (depending on the smallest discrepancy) and $F(\log(1+S_{\text{max}}))=1$.
With the the CDF defined, we compute the AULTC as the area under $F(x)$ from $x=0$ to $x=\log(1+S_{\text{max}})$, normalized by $\log(1+S_{\text{max}})$:
AULTC $=\frac{1}{\log(1+S_{\text{max}})}\Bigg{[}\Big{(}\sum_{i=1}^{k}(x_{(i)}-x_{(i-1)})\frac{i}{k}\Big{)}+(\log(1+S_{\text{max}})-x_{(k)})\times 1\Bigg{]}$
where $x_{(0)}=0$. With this definition, AULTC = 1 indicates that discrepancies are zero (perfect recovery), resulting in maximum area $\log(1+S_{\text{max}})$, and AULTC = 0 indicates that all discrepancies exceed $S_{\text{max}}$, yielding zero area.
Additional remarks on the necessity of explicit bounds for the log-time CDF and AULTC, as well as the non-convexity of the average log discrepancy, are provided in Appendix References.
2.4 Annotation Quality Review
To add to the quantitative assessment of annotations at the event-time level, we also verified the annotation quality at the case-report level in our annotation quality review. To inspect the quality of the annotations, we had our physician annotator review the annotations of the Sepsis-10 dataset that the LLM models produced. The annotator was asked to rank-order the annotations (as well as the manual annotation file), with rank 1 being the best, and rank 7 being the worst. The annotator also placed the annotations into quality levels of Excellent, Good, Acceptable, and Poor. The guideline for placing the annotations into each quality level:
Excellent: High quality extraction. May contain up to 1 clear error in extraction or timing; extracts all remaining important findings; ordering and timing are plausible.
Good: Extracts the important findings, almost always gets the ordering correct, and greater than three-quarters of timings are plausible.
Acceptable: Extracts the important findings, mostly gets the ordering correct, timing is reasonable for more than half of the timing groups.
Poor: Fails to meet criteria for Acceptable.
Annotations of the same quality level were assigned their lower rank (closer to 1). After placing in approximate relative rank, the annotator was asked to pairwise compare all adjacent annotations to ensure relative rankings and ties. The annotator was asked to document pertinent findings and errors in free text.
[…]
Comparison of event match rate (Event), median concordance ($c$), median absolute error (MAE), and area under the log-time curve (AULTC) across LLMs for i2m4, sepsis-10, and sepsis-100 datasets. i2m4 sepsis-10 sepsis-100 Model Event c MAE AULTC Event c MAE AULTC Event c MAE AULTC O1-preview 0.76 0.932 22 0.796 DeepSeek R1 0.35 0.702 16 0.751 0.79 0.876 12 0.772 0.91 0.875 2 0.801 L3.3 70B 0.39 0.730 27 0.721 0.75 0.932 24 0.759 0.78 0.894 4 0.791 L3.1 70B 0.30 0.688 21 0.738 0.75 0.915 24 0.741 0.76 0.891 5 0.790 Mixtral 8x7B 0.23 0.731 27 0.726 0.68 0.914 72 0.570 0.71 0.860 24 0.737 L3.1 8B 0.34 0.707 61 0.682 0.69 0.801 376 0.615 0.73 0.78 48 0.705 BERT 5-CV 0.704 34 0.520
3 Results
Our analysis resulted in T2S2, a new open-access textual time series corpus for Sepsis-3, comprising 2,139 case reports. Demographically, the corpus is 58% male, 42% female, with ages ranging from 0 to 111 and a mean age of 49 (IQR: 32-65). The sepsis-10 manual annotations are of 10 files: PMC3075162, PMC4778217, PMC6238558, PMC7576661, PMC9136437, PMC9355071, PMC9552858, PMC10370906, PMC10556716, and PMC10629858, resulting in 801 reference events. An example annotation set (manual, O1-preview, DeepSeek R1, Llama 3.3, Mixtral 8x7B, and Llama 3.1 8B) for the last case is shown in Appendix .1.
3.1 Phenotyping Results
For phenotyping the i2m4 dataset for sepsis, we found just two case reports of patients with sepsis. The QwQ-32B-Preview response had perfect accuracy (100%), whereas the Llama-3.1-8B-Instruct accuracy was 13/20 (65%). For the sepsis-100 dataset, where QwQ-32B-Preview outputs were labels, the relatively faster models GPT-4o and Llama-3.1-8B-Instruct agreed 84% and 83% of the time, respectively (Figure 1, bottom right). Although the QwQ-32B-Preview responses were repetitive and verbose when making a Sepsis-3 diagnosis, they were visibly based on external foundation knowledge not provided in the prompt, e.g., through use of SOFA rather than qSOFA for phenotyping.
3.2 Annotation Results
The results in Table 2.4 show the strong performance of O1-preview, DeepSeek R1 IQ1, and Llama-3.3-70B-Instruct across all three datasets. In i2m4, all match rates are low (BERT 5-CV is omitted since the method provides the event list [9]), highlighting the difficulty of matching short text spans corresponding to “concepts” rather than chaining them at the level of a clinical finding. Among the matches, Llama 3.3 and DeepSeek R1 performed best for time predictions (trading off event matching with temporal performance), and substantially outperforming the baseline BERT model, particularly with respect to time discrepancy. However, the overall results on i2m4 were considerably lower than in the PMOA datasets, including in terms of time ordering and time discrepancy. In contrast to concordances of 0.707 in i2m4, in sepsis-10, O1-preview and Llama 3.3 were excellent time-ordering annotators, both with concordances of 0.932.
[…]
Comparison of event match rate (Event), median concordance ($c$), median absolute error (MAE), and area under the log-time curve (AULTC) across Llama-3.3-70B-Instruct prompt variants: no role-playing (no role), no conjunction expansion (no expansion), zero-shot, interval, and interval with i2b2 event typing (Interval + Type), for our 3 datasets. i2m4 sepsis-10 sepsis-100 Model Event c MAE AULTC Event c MAE AULTC Event c MAE AULTC L3.3 70B 0.40 0.707 27 0.721 0.75 0.932 24 0.759 0.78 0.894 4 0.791 Ablations No role 0.37 0.670 24 0.738 0.75 0.920 24 0.743 0.74 0.887 4 0.789 No expansion 0.37 0.659 23 0.723 0.73 0.906 24 0.753 0.78 0.886 3 0.793 Zero-shot 0.27 0.632 17 0.741 0.56 0.748 24 0.746 0.69 0.803 6 0.795 Augments Interval 0.41 0.685 15 0.713 0.77 0.921 48 0.725 0.78 0.887 5 0.791 Interval+Type 0.38 0.653 13 0.695 0.66 0.918 24 0.736 0.76 0.867 6 0.776
To examine the results in greater detail, we show figures of the cumulative distribution function of event matches and concordance boxplots (Figure 2) for the sepsis-10 dataset. When the function reaches its maximum at the maximum cosine distance, it may be less than 1 because the label events have no events left to match, i.e., the reference method recovered an event the prediction method did not. The event match rates in Table 2.4 correspond to the match rate at the cosine distance threshold in Figure 2. One could adjust the cosine distance threshold, for example, by increasing it to increase matches, at the cost of false positives, which degrade temporal performance characteristics. The concordance boxplots (Figure 2) show strong performance among the large LLMs, whereas the Llama 3.1 8B and Mixtral 8x7B model performances are much lower—a consistent finding throughout our analysis.
To better visualize the time discrepancies, we plot cumulative time discrepancy plots (Figure 2) that show, for example, that 44% of the O1-preview annotations are exact time matches and that 80% are within 1 week of the label timestamp. Breaking it down further (Figure 2, right), we can see that among events occurring within 1 day of presentation (but more than 1 hour), 90% of O1-preview annotations are within a 1-week time error. This drops to 60% when considering events up to a year from presentation. Problematically, for labeled events more than 1 year from presentation, only O1-preview and DeepSeek R1 ever identify the label time to within 1 year exclusive (Figure 2, right).
The sensitivity analysis to our chosen prompt shows the relative value of various elements for the Llama-3.3-70B-Instruct model (Table 3.2). Our original prompt appears to be the strong performer overall, showcasing the value in particular of few-shot prompting, but also the value of the technical specification of how to treat conjunctions and the modest benefit of role-playing. Surprisingly, when comparing the main L3.3 70B model responses to the augmentations responses, it appears that requesting timestamp intervals (Interval) modestly increases the event match rate. However, the temporal performance characteristic degrades, suggesting a possible request overload on the model or an increase in false positive event matches. The performance of the typing method (Interval + Type) is further degraded. Detailed assessments for the sepsis-10 ablations and the i2m4 dataset, analogous to Figure 2, are shown in Appendix Reconstructing Sepsis Trajectories from Clinical Case Reports using LLMs: the Textual Time Series Corpus for Sepsis. To better understand the performance differences across datasets, we conducted an error analysis between sepsis-10 and i2m4, which is presented in the Appendix .5. This analysis provides additional insight into the factors contributing to variations in model performance for event match rate and c-index across datasets.
Figure 2: Event match CDF (left), concordance box-plots (middle left), time discrepancy from the manual annotation timestamps among matched events, overall (middle right), and disaggregated by clinician annotator timestamp (time from presentation, right) for the sepsis-10 dataset.
3.3 Annotation Quality Results
Inspecting the clinical quality of the annotations, we report the mean rank, top-1 fraction, top-1-LLM fraction, and quality level in Table 4. There is a clear performance ordering, with O1-preview annotations being a close second to the manual annotations, followed by DeepSeek R1 IQ1 and Llama 3.3. While the Llama 3.3 annotations are found to be acceptable a high percentage of the time (90%), the other Llama and Mistral approaches fail to achieve Acceptable performance in excess of 50%. Comparing the DeepSeek R1 IQ1 and Llama 3.3 annotations, the former has more variable performance (only 8 of 10 with Acceptable annotations), but maintains both higher match rate and AULTC (Table 2.4 and Figure 2 (left and middle right)). Our clinical reviewer commented that the Llama 3.3 discrepancies for large time horizons, i.e., years, were large, which is corroborated in Figure 2 (right, bottom subpanels).
The annotation review feedback noted a wide variety of errors; here we present several errors modes identified during quality review.
Events with duration. Example: “10-kg weight loss” from “he was presented to the hospital with complaints of abdominal distension, constipation, vomiting, and a 10-kg weight loss.” This example from case report PMC10629858 references a gradual event that has been ongoing for 2 months. By the annotation guideline for selecting the start time of events with duration, the time -168 (what O1-preview annotates) is correct. However, the event can also be considered as “history of 10-kg weight loss” via the “history of” guideline with time 0 (what L3.3 70B annotates). Note, this example could be resolved by interval representations, but we see the annotation quality (particularly the AULTC) degrade when requesting additional fields (Table 3.2 and Figure .3).
Events requiring temporal reasoning or composition. Example: “passed away” from “he was re-admitted to the medical ICU for severe sepsis and multiorgan failure and passed away around 6 months after his initial diagnosis with NHL, despite maintaining a remission status.” This requires a two-step inference, (1) timing of NHL diagnosis, and (2) adding “around 6 months” to that time. While this is a simple example, more complicated chains of relative timings were noted, and possibly relative to other events whose timings are also uncertain. As black-box methods, there is no explicit probabilistic specification for these chains, and empirically we see variable responses. For this case, the manual annotation is: 4383 (about 6 months), O1-preview gives 4368 (about 6 months), DeepSeek R1 gives 4032 (about 6 months), and L3.3 70B gives: 1440 (about 2 months; compared to other events the temporal ordering is correct, but the relative timing is off by months).
4 Discussion and Conclusions
We developed an analytical pipeline for processing free text biomedical corpora to create textual time series for Sepsis-3 using LLM and showed strong performance characteristics, including good match rates, high concordance and low time discrepancies. We plan to release the full cohort annotations upon acceptance, providing a novel Sepsis-3 cohort with rich textual information, annotated with timestamps, which may be useful for downstream tasks such as risk forecasting, causal reasoning, and disease trajectory characterization. In PMOA samples sepsis-10 and sepsis-100, performance levels were strong with respect to clinician annotations (sepsis-10) and inter-LLM agreement (sepsis-100).
[…]
Sepsis-10 annotation quality review Annotator Mean Rank Top-1 Top-1-LLM Excellent $\geq$ Good $\geq$ Acceptable Manual 1.2 0.9 1 1 1 O1-preview 1.5 0.6 0.9 0.9 1 1 DeepSeek R1 IQ1 3.5 0.1 0.2 0.4 0.5 0.8 L3.3 70B 3.8 0 0 0 0.2 0.9 L3.1 70B 4.8 0 0 0 0.2 0.5 Mixtral 8x7B 5.4 0 0 0 0.1 0.2 L3.1 8B 6.5 0 0 0 0 0
There are several limitations to our work. First, our product, the textual time series corpus for sepsis, is a clinical research informatics cohort, and possesses its characteristic limitations. It should not be considered a representative sample, since case report corpora possess biases, including but not limited to publication biases, rare finding over-representation, and revisionism. Second, while we analyzed multiple data sources, we are limited by the number of clinician annotations that are challenging and expensive to obtain. Third, the substantial use of LLMs in the analytic pipeline can make error analysis more difficult, as their errors can be subtle and inaccurate.
Regarding generalizability to other medical conditions besides sepsis, our methodology could be extended in a straightforward manner. Our motivation for choosing sepsis as the disease focus was to expand the analysis from the i2m4 dataset to a similar population in the PMOA corpus and maintain a well-defined phenotype. Sepsis is a sensible choice as it is one of the most prevalent and important diagnoses in critical care medicine. However, there are several concerns when generalizing to other conditions. For one, most other conditions are less acute in nature, meaning that the time distributions may be wider. Also, the amount and type of information in such case reports may vary. Despite this, one of the benefits of the AULTC (compared to linear/squared discrepancies) when mixing intensive care and chronic disease cases is that the chronic disease discrepancies at wider times scales do not dominate the performance measure (because of the time discrepancy log scaling).
Another generalizability concern is that for some conditions, less temporal information will be present. This appears to be more acute in imaging studies, where the focus of the case is on the image content at a single snapshot in time. For these types of studies, the longitudinal annotations become less meaningful compared to the image’s information content.
Future work could further characterize the findings, with additional or improved augmentations to the current system. Although our augmentation strategies did not improve overall system performance (Table 3.2), alternative prompting and or fine-tuning frameworks could further improve the annotations in terms of quality and representation. Our hope is that this corpus, or a future version of it, will be used for downstream applications in sepsis, such as causal modeling, forecasting, and multimodal alignment. As a community, we must use the best data available regardless of its format, and for temporal tasks, this temporal, textual approach offers a complementary corpus to those in frequent use. Further developments in temporal extraction could improve upon our pipeline and could be tailored to one or more of the applications above, each of which place different emphasis on what constitutes a high-quality clinical time series. For example, time-ordering is a necessary condition for causality, whereas the precise timestamp may be of high importance in forecasting and alignment. Finally, while our clinical finding and timestamp pairs attempt to capture the true occurrence and timing of each event, another approach focusing on the timing of perceived occurrences could better uncover how actors process and affect their environments, leading to better characterization for decision analysis, multi-agent systems, and resource allocation. Multimodal integrations could help account for perception differences and design collection strategies to assist early detection of sepsis and sepsis progression.
5 Acknowledgments
This research was supported in part by the Division of Intramural Research (DIR) of the National Library of Medicine (NLM), National Institutes of Health. This work utilized the computational resources of the NIH HPC Biowulf cluster.